OpenStudy (piercetheveil47):

Multiply (2.0 times 10^−4) times (3.1 times 10^−20). Express the answer in scientific notation. 6.2 times 10^−80 6.2 times 10^−24 6.2 times 10^24 6.2 times 10^80

OpenStudy (anonymous):

\[(2.0 \times 10^{-4}) \times (3.1 \times 10^{-20})\] First since you are working with multiples of 10 you can multiply the 2.0 and 3.1 to find the 6.2 and then apply the laws of exponents where you have to numbers with the same base. \[A^{X} \times A^{Y} = A^{X+Y}\]

OpenStudy (piercetheveil47):

is the answer b??

OpenStudy (anonymous):

In your case you have two negative numbers which would add up to a larger negative number.

OpenStudy (anonymous):

Correct

OpenStudy (piercetheveil47):

thank you!

OpenStudy (anonymous):

No problem